OpenStudy (anonymous):

Find the x value for point F such that DF and EF form a 1:3 ratio. D(-3, -3), E(2, 4)

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

honestly me too. @KamiBug @StudyGurl14

OpenStudy (anonymous):

They should be able to help

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yep

OpenStudy (studygurl14):

Okay hold on a sec.

OpenStudy (anonymous):

ok

OpenStudy (studygurl14):

you have to find a point on the graph that will make the ratio of the lengths of sides DF and EF to be 1:3

OpenStudy (anonymous):

yes i understood that. do you think it is -1.75

OpenStudy (anonymous):

3+2=5 distance 5*1/4=1.25 1.25-3=-1.75

OpenStudy (anonymous):

taht the only choive that can be made

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

look at the grath. it looks like this

OpenStudy (studygurl14):

what is the coordinate @ttb123456789 that's the only way i can check...

OpenStudy (anonymous):

we should find only the x value.

OpenStudy (anonymous):

@StudyGurl14 are you there

OpenStudy (anonymous):

i'm pressing sumbit. Are you sure that i'm correct?

OpenStudy (studygurl14):

I don't know...i think so.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@KamiBug

OpenStudy (kamibug):

Yeah I think that must be pretty close. I'd estimate the x value is near -1.6. So go for it! ^.^

OpenStudy (anonymous):

Geart Thank you guy